OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (2, 2) and (4, 8)? (y - 2) = 3(x - 2) (y + 2) = 3(x + 2) (y - 8) = -3(x - 4) (y + 8) = -3(x + 4)

OpenStudy (princeharryyy):

its the first one @andy11

OpenStudy (princeharryyy):

plug in y-y1 = m(x-x1) ; m = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

you sure?

OpenStudy (princeharryyy):

yup i am @andy11 100% sure

OpenStudy (anonymous):

ok thank you! do you think you can you help with 1 more ?

OpenStudy (anonymous):

What is the slope of the line which passes through (4, 7) and (2, 3)? 2 -2 3 -5

OpenStudy (princeharryyy):

(3-7)/(2-4) => -4/-2 =2

OpenStudy (princeharryyy):

@andy11

OpenStudy (anonymous):

thank you so much! do you know anything about rate of change?

OpenStudy (princeharryyy):

yup I do

OpenStudy (anonymous):

Use the table below to answer this question: x y -1 7 3 3 5 1 Find the average rate of change for the given function from x = -1 to x = 5. -6 -1 1 6

OpenStudy (princeharryyy):

(1-7)/(5-(-1)) = -1 ans

OpenStudy (princeharryyy):

@andy11 i was having my food so a bit late

OpenStudy (princeharryyy):

u there ??/ @andy11

OpenStudy (anonymous):

oh ok thats fine thank you again

OpenStudy (princeharryyy):

u r welcome @andy11